OpenStudy (anonymous):

Which of the following points lies on the graph of the inverse of f(x) = 6^(x + 1)? I know the answer is (1, -1). But how do you find out ? Because when I graphed it, I got (-1, 0). So I'm not sure what I did differently ? When I put it into my calculator I went to the y= button & typed in log(6)(x+1) because the inverse of f(x) = 6^(x + 1) is f(x) = log(6)(x+1). Can someone help me find my mistake ?

OpenStudy (anonymous):

Or could the inverse be f(x)=log(6)(x-1) ? But even after graphing that I got (1,0). So I did something wrong somewhere ..

OpenStudy (anonymous):

You made the mistake while U find the inverse function

OpenStudy (anonymous):

It should be\[f^{-1}(x)=(\log_{6}x)-1\]

OpenStudy (anonymous):

But wouldn't the inverse function be f(x)=log(6)(x-1) ? Because looking at the other problems I had, that would make sense I thought.. So if it isn't that, what would it be ?

OpenStudy (anonymous):

f^-1(x) = (logx)/7 - 1

OpenStudy (anonymous):

Well in the next problem on my work, it said find the inverse of f(x) = 3^(x - 1) & the correct answer was log(3)(x+1).

OpenStudy (anonymous):

So wouldn't the inverse of f(x)=6^(x+1) be f(x)=log(6)(x-1)?

OpenStudy (anonymous):

yep it is what i wright before

OpenStudy (anonymous):

By the way did U know how to find the inverse of the function?

OpenStudy (anonymous):

Yes, I know how to find the inverse. It's just the signs that threw me off this time. & Since we have the right inverse, how do we graph it ? Because the first way I graphed it was apparently wrong...

OpenStudy (anonymous):

First find the graph of \[\log_{6}x\]and apply vertical shifting rule to find out the graph of\[(\log_{6}x)-1\]

OpenStudy (anonymous):

@bhaskarbabu

OpenStudy (anonymous):

I have no idea what the vertical shifting rule is

OpenStudy (anonymous):

Sorry ! I've never had to use it before...

OpenStudy (anonymous):

Shift the first graph vertically

OpenStudy (anonymous):

How ?

OpenStudy (anonymous):



OpenStudy (anonymous):